# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ 6 Training with Synthetic Data
Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with straightforward training methods, thereby addressing the challenge of insufficient training data. We first show that just joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1). We then go a step further and show that a simple Curriculum Learning method can further boost performance (Section 6.2). The synthetic data together with the original BAP (augmented) data now constitutes the BAP v2 training dataset, completing the BAP v2 framework. We also illustrate the impact of the synthetic data and training methodologies on training a simple LLM (BERT) and transformer-based model (Section 6.3).
A note on experimental design choices
For the GRU-based models in Sections 6.1 and 6.2, we maintain the same architecture and hyperparameters as the baseline model (Section 2.6), along with the same amount of data augmentation in the BAP training data as in our prior work (Section 2.5). Fine-tuning these aspects for different training methods is beyond the scope of our objectives here and left for future work to further optimize performance. Similar choices made for the BERT-based models are discussed later in Section 6.3.
@@
Hereafter, we refer to the random structure-based dialogs (Section 5.2) as $D_{\mathrm{r}}$, blocks-based dialogs for shape-based structures (Section 5.3.2) as $D_{\mathrm{bs}}$, and shape-based dialogs for shape-based structures (Section 5.3.3) as $D_{\mathrm{ss}}$. The Minecraft-based original BAP data is denoted as $D_{\mathrm{mc}}$. Each dataset includes train, test, and val splits, with the $D_{\mathrm{mc}}$ train set also incorporating augmented data from Section 2.5. The baseline model (Section 2.6), trained on $D_{\mathrm{mc}}$ alone, is denoted as $M_{\mathrm{mc}}$.
6.1 Joint training on aggregated data
The most straightforward approach to use the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) is to just combine them all together along with the original $D_{\mathrm{mc}}$ data and retrain the baseline model on this aggregated dataset.
@@ 6.1.1 Experimental Setup
Our setup closely follows the one used for the baseline model $M_{\mathrm{mc}}$ when it was trained on $D_{\mathrm{mc}}$ alone (Section 2.6). We retain the same exact model architecture and hyperparameters as in $M_{\mathrm{mc}}$. Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions. Further details are provided in the appendix (Section 12.3). We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.4) and the test sets of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.  [^6] For fairer F1 on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1). This is because, by design, all EB items in $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations. Results on the full v2 evaluation benchmark are reported for the baseline and best performing models in this work later in Section 7.1.
@@ 6.1.2 Experiments and Results
Table 4 presents the results. We denote the model trained on the aggregated dataset as $M_{\mathrm{agg}}$, and include comparisons to the baseline model ($M_{\mathrm{mc}}$– trained on $D_{\mathrm{mc}}$), and models trained on the three synthetic datasets separately ($M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$– trained on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ respectively). F1 performance is reported (as percentages) on the respective test sets of all four datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$).
$M_{\mathrm{agg}}$ achieves an F1 score of 30.3% on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3%, a notable 3-percentage point improvement. Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these. For a sense of what high performance means on these datasets, we reference the diagonal entries in rows for models $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, representing performance of these models on the same test data distribution as train: 83.9% for $D_{\mathrm{bs}}$, 75.4% for $D_{\mathrm{ss}}$, and 64.4% for $D_{\mathrm{r}}$. In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$. It is also better on the synthetic datasets than on $D_{\mathrm{mc}}$ by large margins (as it should be), unlike $M_{\mathrm{mc}}$. Lastly, the models trained solely on synthetic datasets perform poorly on $D_{\mathrm{mc}}$, as does $M_{\mathrm{mc}}$ on synthetic data, indicating that joint training on the aggregated dataset is essential for robust performance across both real and synthetic data.
@@
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
@@
$M_{\mathrm{agg}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
Table 4: Models trained on datasets separately ($M_{\mathrm{mc}}$, $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$) and together ($M_{\mathrm{agg}}$)
6.2 Curriculum Learning (CL)
Another key observation from Table 4 is that performance of the $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, and $M_{\mathrm{mc}}$ models decreases in the order $M_{\mathrm{bs}}$ > $M_{\mathrm{ss}}$ > $M_{\mathrm{r}}$ > $M_{\mathrm{mc}}$, suggesting that the complexity/difficulty of examples increases in the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$, with the most complex $D_{\mathrm{mc}}$ examples naturally appearing at the end. This observation raises the possibility of using this ordering to improve over the naive joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$. Curriculum Learning (CL) is a natural next step to explore in this direction. CL is a training strategy that sequences training data from simpler to more complex examples, similar to human learning. Introduced by Bengio et al. (2009), CL begins with easier examples, gradually increasing in difficulty. This progression allows the model to first learn basic patterns and representations, building a foundation before tackling more challenging cases, which can enhance generalization and improve performance on complex tasks.
6.2.1 CL Mechanics
Difficulty Metrics
A key prerequisite in CL is establishing difficulty metric(s) that quantify how difficult a training example is for a model to learn, allowing the data to be sorted accordingly during training. In this work, we explore three straightforward, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another). The first metric is the dataset $D$ an example belongs to ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$). As mentioned before, the datasets can be ordered as $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$ in terms of difficulty. Additionally, we consider the lengths of the action (output) and input sequences: the action (output) sequence length $L_{A}$ reflects the number of actions in the gold action sequence, while the input sequence length $L_{I}$ measures the number of tokens in the game history fed to the model. These are also intuitive as longer sequences are harder to generate as output and process as input.
CL Algorithm
We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al. (2020). Details of the method are provided in the appendix (Section 12.1).
@@ 6.2.2 Experiments and Results
Overall experimental setup remains similar to the one used for joint training (Section 6.1.1). We experiment with each of the three difficulty metrics individually and the four possible combinations of them, making a total of seven metrics – $D$, $L_{A}$, $L_{I}$, $(D,L_{A})$, $(D,L_{I})$, $(L_{I},L_{A})$, and $(D,L_{I},L_{A})$ (ordering within a tuple denotes the sorting order). We also experiment with the CL algorithm’s parameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30. (This controls the number of buckets/difficulty levels – see appendix (Section 12.1).) This results in a grid of $7\times 6=42$ experiments.
Table 5 presents the best model, $M_{\mathrm{agg+cl}}$, selected based on $D_{\mathrm{mc}}$ performance. For comparison, we also include the baseline $M_{\mathrm{mc}}$, and the aforementioned $M_{\mathrm{agg}}$ model (Table 4). On $D_{\mathrm{mc}}$, $M_{\mathrm{agg+cl}}$ achieves an F1 of 33.0%, a notable 2.7-point improvement over $M_{\mathrm{agg}}$, and an even more significant 5.7-point improvement over the baseline $M_{\mathrm{mc}}$. It also continues to be highly performant on the synthetic data similar to $M_{\mathrm{agg}}$. Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over naive joint training. Hereafter, we refer to $M_{\mathrm{agg+cl}}$ as the best model.
Model Train Data Test Data
$D_{\mathrm{bs}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
$M_{\mathrm{mc}}$ $D_{\mathrm{mc}}$ 35.2 19.1 12.4 27.3
$M_{\mathrm{agg}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
$M_{\mathrm{agg+cl}}$ $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 84.0 68.8 60.3 33.0
Table 5: Curriculum learning ($M_{\mathrm{agg+cl}}$) compared to models $M_{\mathrm{mc}}$ and $M_{\mathrm{agg}}$
6.3 Transformers and LLMs (BERT)
As noted in Section 1, modeling is not the primary focus of this work. However, solely to test the robustness of our synthetic data and training methodologies and show their applicability to relatively contemporary model architectures, we experiment with a very simple alternative version of the baseline model (Section 2.6) that incorporates a Transformer architecture (Vaswani et al., 2017) and an LLM (BERT; Devlin et al. (2019)). Specifically, we replace the GRU-based encoder (and GloVe embeddings) with BERT and substitute the GRU decoder with a Transformer, keeping the rest of the architecture unchanged (Figure 4). (A projection layer is added atop BERT for dimensionality reduction.) The simplicity also allows us to assess whether/to what extent simpler LLMs like BERT and vanilla Transformers alone can already effectively handle the BAP task. Future work can build on the insights we obtain to explore more sophisticated LLMs and architectures while leveraging the BAP v2 framework.
Model details
To control the representation dimensionality and complexity of the decoder transformer, we add a feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be). The BERT encoder and transformer decoder are connected in standard encoder-decoder fashion but via the projection layer. BERT’s token embeddings are compressed by the projection layer, and the transformer decoder then attends to these representations at each decoding time-step, and the representation dimensionality of the decoder matches the output size of the projection layer. This also matches the embedding dimensionality for action vectors fed into the decoder. Positional embeddings are also used for the decoder input. The output of the decoder’s last layer conditions the CNN-based action predictor (replacing the GRU hidden state from the baseline model).
6.3.1 Experiments and Results
We conduct similar experiments with this model like with the GRU-based model, i.e., training on $D_{\mathrm{mc}}$ only, joint training on the aggregated dataset, and CL. The experimental setup and the CL experiments are the same as for the GRU scenario (Sections 6.1.1 and 6.2.2). Further details specific to the BERT scenario are provided in the appendix (Section 12.4). We tested various sizes of BERT models (Turc et al., 2019), and keeping BERT frozen versus finetuning it. We found that frozen $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed the best. In the transformer decoder, five layers yielded optimal results, and a projection layer output size of 256 proved most effective. Other hyperparameters are provided in the appendix (Section 12.4). For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g. those of the CNNs). Results are shown in Table 6. (Note: We report only the best-performing BERT-based model, which was from the frozen scenario as mentioned earlier. However, similar to the results for the frozen case discussed below, synthetic data, joint training, and CL also improved model performance in the finetuning scenario, validating our approach.)

### v3

@@ 6 Revisiting BAP Modeling: Training with Synthetic Data
    Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with straightforward training methods, thereby addressing the challenge of insufficient training data, and thus validating one of the motivations behind our synthetic data. We demonstrate this for both non-LLM- and LLM-based models, highlighting the robustness of our approach. The synthetic data together with the original BAP training data now constitutes the BAP v2 training data in the BAP v2 framework.
@@
    Recall that we refer to the synthetic datasets as $D_{\mathrm{r}}$, $D_{\mathrm{sb}}$ and $D_{\mathrm{ss}}$. The Minecraft-based original BAP data is denoted as $D_{\mathrm{mc}}$. Each dataset includes train, test, and val splits. The GRU-based baseline model (Section 2.6), trained on $D_{\mathrm{mc}}$ alone, is denoted as $M_{\mathrm{mc}}$.
    6.1 The GRU-based models
    The most straightforward approach to use the synthetic datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) is to just combine them all together along with the original $D_{\mathrm{mc}}$ data and retrain the GRU-based baseline model on this aggregated dataset.
@@ 6.1.1 Experimental Setup
    Our setup closely follows the one used for the baseline model $M_{\mathrm{mc}}$ when it was trained on $D_{\mathrm{mc}}$ alone (Section 2.6). Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions. Further details are provided in the appendix (Section 11.2). We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.3) and the test sets of $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.  [^9] For fairer F1 on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1). This is because, by design, all EB items in $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations.
@@ 6.1.2 Experiments and Results
    Table 4 report overall F1 performance on the respective test sets of all four datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$) for all variants of our GRU models. We denote the model trained on the aggregated dataset as $M_{\mathrm{agg}}$, and include comparisons to the baseline model ($M_{\mathrm{mc}}$, trained on $D_{\mathrm{mc}}$), and models trained on the three synthetic datasets separately ($M_{\mathrm{sb}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, trained on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ respectively).
    $M_{\mathrm{agg}}$ achieves an F1 score of 30.3 on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3, a notable 3- point improvement. Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these (83.4 vs. 35.2 on $D_{\mathrm{sb}}$, 68.8 vs. 19.1 on $D_{\mathrm{ss}}$, and 63.6 vs. 12.4 on $D_{\mathrm{r}}$). It is also informative to compare both the $M_{\mathrm{mc}}$ and $M_{\mathrm{agg}}$ numbers with the corresponding diagonal entries in the $M_{\mathrm{sb}}$, $M_{\mathrm{ss}}$ and $M_{\mathrm{r}}$ rows, which report the performance of these models when trained only on these datasets: 83.9 for $D_{\mathrm{sb}}$, 75.4 for $D_{\mathrm{ss}}$, and 64.4 for $D_{\mathrm{r}}$ with the baseline $M_{\mathrm{mc}}$ numbers. $M_{\mathrm{mc}}$ clearly underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$. Lastly, the models trained solely on synthetic datasets perform poorly on $D_{\mathrm{mc}}$, as does $M_{\mathrm{mc}}$ on synthetic data, indicating that joint training on the aggregated dataset is essential for robust performance across both real and synthetic data.
@@
    $D_{\mathrm{sb}}$ $D_{\mathrm{ss}}$ $D_{\mathrm{r}}$ $D_{\mathrm{mc}}$
@@
    $M_{\mathrm{agg}}$ $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$ 83.4 68.8 63.6 30.3
    Table 4: Models trained on datasets separately ($M_{\mathrm{mc}}$, $M_{\mathrm{sb}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$) and together ($M_{\mathrm{agg}}$)
    6.2 LLMs (Llama)
    To demonstrate the robustness of our synthetic data and its applicability to contemporary Large Language Models (LLMs), we conduct experiments with Llama (Dubey et al., 2024). Specifically, we consider the Nebula model recently proposed in Chaturvedi, Thompson, and Asher (2024) that established a new SOTA on the BAP task and aim to further improve upon it. We refer to it as our LLM-based baseline model. It is a Llama-3-8B model that takes as context all conversation and action sequences up to the action sequence $A$ to predict $A$ in a text-to-text fashion. It is finetuned on the Minecraft data $D_{\mathrm{mc}}$ using QLoRA (Dettmers et al., 2023).
@@ 6.2.1 Experiments and Results
    We conduct experiments analogous to those performed with the GRU-based models, along with additional experiments on the input textual representation provided to the LLM. The results are detailed in Table 6. All models involve finetuning Llama-3-8B using QLoRA.
    LLM-based baseline:
    First, we reproduce the Nebula LLM-based baseline by fine-tuning Llama-3-8B on $D_{\mathrm{mc}}$ but this time evaluating it on $D_{\mathrm{mc}}$ using the v2 evaluation as well as on the synthetic data. As shown in the first row of Table 6, this model ($M_{\mathrm{mc}}^{\prime}$) achieves an F1 score of 47.0 on $D_{\mathrm{mc}}$, significantly better than its GRU-based counterpart $M_{\mathrm{mc}}$ (27.3). However, similar to $M_{\mathrm{mc}}$, it underperforms on the easier synthetic data ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), despite being trained on the far more complex $D_{\mathrm{mc}}$, albeit its performance is better than that of $M_{\mathrm{mc}}$. Thus, this counterintuitive behavior of the GRU-based model still persists even with the far more advanced LLM-based model.
    Joint training on aggregated data
    Next, we perform joint training on the aggregated dataset ($D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$) to produce model $M_{\mathrm{agg}}^{\prime}$. This yields substantial improvements on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, as expected, but only a marginal gain on $D_{\mathrm{mc}}$ (47.8 compared to 47.0. Furthermore, its performance on the synthetic datasets still lags behind its GRU-based counterpart, $M_{\mathrm{agg}}$.
    Input Representation: NN
    Predict the action sequence (AS) for the Minecraft excerpt:
    <Architect> place one in the center of one of the lines near the edge
    place yellow 5 1 0
    <Architect> good; repeat to make a V
    place yellow 3 1 -1
    place yellow 3 1 1
    <Architect> still diagonal from the previous squares
    <Builder> extended?
    <Architect> not really; the first step is orange, but it is in line with the yellow blocks
    pick 3 2 1
    pick 3 1 1
    pick 2 2 2
    <Architect> now one orange on top of teh second orange
    place orange 1 2 2
    <Architect> now delete the second orange
    pick 1 1 2
    <Architect> good repeat to make a third step
    ### AS:
    Input Representation: $N+\textsc{Pos}_{\textbf{B}}$ (appends builder’s position and orientation to the input above)
    Builder’s current position is 0.2 1.0 -0.4 and orientation (yaw) is -6.5 degrees
    ### AS:
    Input Representation: $N+\textsc{Pos}_{\textbf{B}}+S$ (appends current built structure to the input above)
    Current built structure is:
    yellow 5 1 0
    yellow 4 1 1
    yellow 4 1 -1
    yellow 3 1 2
    yellow 3 1 -2
    orange 2 1 2
    orange 1 2 2
    ### AS:
    Output (Common to all inputs)
    place orange 1 3 2
    place orange 0 3 2
    pick 1 3 2
    Table 5: The three different input representations ($N$, $N+\textsc{Pos}_{\textbf{B}}$, $N+\textsc{Pos}_{\textbf{B}}+S$) and the (common) ground truth output. Ellipses (…) denote truncation for readability. $N$ is from Chaturvedi, Thompson, and Asher (2024), and $N+\textsc{Pos}_{\textbf{B}}$, $N+\textsc{Pos}_{\textbf{B}}+S$ are ours.
    Richer input representations
    We hypothesize that the lower-than-expected results of $M_{\mathrm{agg}}^{\prime}$ stems from the input representation, which lacks a critical component of the game context available to the GRU models – the Builder’s position and orientation, $\textsc{Pos}_{\textbf{B}}$, which is essential for correctly interpreting spatial relations (Section 2.4). The GRU models incorporate $\textsc{Pos}_{\textbf{B}}$ (Jayannavar, Narayan-Chen, and Hockenmaier, 2020), making the comparison with the LLM models uneven. We therefore enrich the input text by appending information about $\textsc{Pos}_{\textbf{B}}$, as illustrated in Table 5 (representation $N+\textsc{Pos}_{\textbf{B}}$). We denote the original input representation used in Nebula as $N$ and show that at the very top of the table. (The ground truth output representation is shown at the very bottom.)
    Training on the aggregated data with this richer input yields model $M_{\mathrm{agg+posb}}^{\prime}$, which shows significant improvement on $D_{\mathrm{mc}}$ and even more on the synthetic data, in comparison to $M_{\mathrm{agg}}^{\prime}$. It achieves F1 scores of 93.4 on $D_{\mathrm{sb}}$, 84.9 on $D_{\mathrm{ss}}$, and 82.3 on $D_{\mathrm{r}}$, while also boosting the $D_{\mathrm{mc}}$ score to 52.5, and now also overtakes the performance of the GRU models on all datasets, validating out hypothesis that suitable input representations are essential for this task.
    However, to verify that this gain is not from the representation alone, we trained a model ($M_{\mathrm{mc+posb}}^{\prime}$) on just $D_{\mathrm{mc}}$ but with the new representation. This results in an F1 score of 47.6 on $D_{\mathrm{mc}}$, a marginal improvement over the 47.0 from its counterpart $M_{\mathrm{mc}}^{\prime}$ that used the original Nebula representation. This confirms that both the enriched input representation and the synthetic data are necessary to achieve significant performance gains on both synthetic and real Minecraft data; each on their own are not enough. The explicit spatial context in the input allows the model to better leverage the patterns in the synthetic data and transfer that knowledge to the $D_{\mathrm{mc}}$ domain.
    We also note that while our textual representation of $\textsc{Pos}_{\textbf{B}}$ is fairly straightforward, its effect on performance is very substantial. This input only explicitly provides the Builder’s frame of reference, but for the model to correctly ground spatial instructions, it must "infer" a more complex, situated understanding of the 3D world state from this frame. Our results demonstrate that the LLM seems to become, to a significant extent, capable of this complex inference, just through the textual signal in the data and standard training (teacher forcing and cross entropy loss). Crucially, however, this ability due to $\textsc{Pos}_{\textbf{B}}$ only emerges when the model is trained on enough signal from the rich spatial examples in our synthetic data in conjunction with the original $D_{\mathrm{mc}}$.
    A further boost
    Finally, we investigate whether an explicit representation of the current world state/built structure $S$ (like in the GRU-based models) helps further. While the action history contained in the input text already "implicitly" defines $S$, we append an explicit version of it to the input (representation $N+\textsc{Pos}_{\textbf{B}}+S$ in Table 5). Joint training on the aggregated data with this input produces model $M_{\mathrm{agg+posb+s}}^{\prime}$ which further boosts performance on $D_{\mathrm{mc}}$ by a bit to 53.0 (up from the 52.5 of $M_{\mathrm{agg+posb}}^{\prime}$), while performance on $D_{\mathrm{sb}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ continues to remain strong and comparable to $M_{\mathrm{agg+posb}}^{\prime}$.
    This result suggests that for the longer and more complex dialogues in $D_{\mathrm{mc}}$, where tracking and inferring the world state/built structure $S$ from just the textual action history is more challenging, an explicit representation of $S$ provides a beneficial signal. On the shorter and simpler dialogues in the synthetic datasets, however, this inference is not that challenging, and therefore explicit state information may not be as necessary. This likely explains why adding in this representation only improves performance on $D_{\mathrm{mc}}$. Moreover, given the already strong performance on the synthetic data after including $\textsc{Pos}_{\textbf{B}}$, further gains from additional inputs such as $S$ may anyway be subject to diminishing returns on the synthetic data and the transfer of knowledge to the $D_{\mathrm{mc}}$ domain (unlike for $M_{\mathrm{agg+posb}}^{\prime}$ as discussed above).
    We designate $M_{\mathrm{agg+posb+s}}^{\prime}$ as our best model. As it makes use of richer game context and synthetic data (with each one enabling the other), we name it Llama-CRAFTS (Context Rich And Fine-Tuned On Synthetic Data). This model establishes a new state of the art (SOTA) on the BAP task ($D_{\mathrm{mc}}$), achieving an F1 score of 53.0. This represents a 6 point improvement over the 47.0 F1 score of the previous SOTA, the Nebula (the LLM baseline).
    6.3 Broader Takeaways